User name peter
Log entry time 08:01:36 on December18,2003
Entry number 114569
keyword=Owl Shift Summary
Owl Shift Summary
SL: Peter Monaghan
TO: Alexandre Camsonne
3rd: Luigi Lagamba
00:00 Shift begins with problems with the BCM and Harp scans (see
previous shift summary). Doug Higinbotham is currently working
with MCC to try and solve the problems.
00:30 Doug has returned from MCC to tell us that they can do a
Harp scan - perhaps a software issue with our Harps. The BCM
is not working and needs attention. MCC is having problems
their fast feedback system which they are going to look at
00:40 The only test we can do tonight is check the beam spot on the
Be) target; moved target into position (at least it works!).
MCC sending 5uA of unrastered beam; beam spot shows up fine.
00:45 Checking the raster setting with spot++. Ended run # 1402
(cosmics) and start another run #1403 check the raster spot.
01:25 MCC has tried to set as small a raster as possible - it is
showing a spot size of ~ 1 x 2 mm on Spot++; this translates
into 1.6 x 1.45 (x,y) in 'MCC' units. We cannot make a
smaller raster than this - run coordinator and others can
decide if this is sufficiently small enough for the waterfall
01:30 Moving the 'empty' target (solid 5) into position.
01:35 We have nothing else to do this shift due to the problems we
are having. MCC are going to work on steering the beam.
03:45 MCC have finished their own beam steering studies and have
nothing left which they want to do - Goodnight MCC!
06:45 MCC puts Hall A in Restricted Access.
08:00 Shift ends with nothing done and more work required in the hall today.
1402 Cosmic Run
1403 Cosmic Run; checking raster size.
1404 Cosmic Run; checking raster size.
1405 Cosmic run; checking raster size.
1406 Cosmic run, with RICH in trigger.